OpenStudy (anonymous):

Alebra Urgent!!!: A crew of 30 people can build a cettain road in 60 days. After the tenth day the plans are changed; the company wants the road built in 30, not 60, days. How many more people must be hired?

OpenStudy (anonymous):

They have to hire 60 people. I think...

OpenStudy (anonymous):

thnx, but do you know lik ethe steps to solve it?

OpenStudy (anonymous):

oh kk lol. thnx for trying :)

OpenStudy (amistre64):

how much of the job is done in one day by the 30 people?

OpenStudy (anonymous):

Listen to Amistre...he's smart.

OpenStudy (amistre64):

1 job = 60 days ; we want the amount of job done in 1 day, so divide by 60 1/60 job = 1 day 1/60 job = 30 people ; there are 30 people, and we want to know how much of the job a single person does in 1 day; divide by 30. 1/1800 job = 1 person now, how much of the job is left will give us a good idea of how many people are needed

OpenStudy (amistre64):

30 people for 10 days: 300/1800 = 1/6 of the job done in 10 days now they want the rest of the road built in the next 20 days and we have 5/6 of the job left to complete. p*20/1800 = 5/6 should give us how many people are needed to complete the job in that time frame, p-30 = how many people need to be hired on since we already have 30 people to work on it

OpenStudy (anonymous):

Oh, ok so i did it and i got 75

OpenStudy (anonymous):

so 75-30=45 so we need 45 more workers. Oh nice, thanks!!!